# results on dynamics of bungee set of composite entire functions in the eremenko-lyubich class

### v3

@@ 3. Some results for functions of class $\mathcal{B}$
Now, we will prove some results when these functions indeed have some finite asymptotic values, and to simplify, we will focus on the functions of the Eremenko-Lyubich class $\mathcal{B}$. These functions do not have a Baker domain; therefore, the escaping set does not intersect with the Fatou set. In addition, they do not have a finite intersection of the bungee set with the wandering domain because the wandering domain of such functions does not have (sub)sequences that tend to $\infty$.
Therefore, the Julia set for functions of class $\mathcal{B}$ comprises of $RP(f),BU(f)$ and $I(f)$; that is, $J(f)=BU(f)\cup I(f)\cup RP(f)$. For commuting functions $f$ and $g$, we have $J(f\circ g)\supset J(f)\cup J(g)$ [10, 11]. However, when these are commuting functions of class $\mathcal{B}$ we have, $J(f)=J(g)=J(f\circ g)$ [10]. We also know that $g(J(f))\subset J(f)$ and $\overline{RP(f)}=J(f),\overline{I(f)}=J(f)$ [4, 6].
@@
The bungee set for functions of class $\mathcal{B}$ does not intersect with Fatou set.
@@
[…]
This also means that $BU(f)\subset J(f)$ only for functions of class $\mathcal{B}.$ Also, it is easy to see that $BU(f)$ is completely invariant by the complete invariance of $I(f)$ and $K(f)$. As a result, by the minimality property of Julia set, $\overline{BU(f)}=J(f)$.
@@
The bungee set for functions of class $\mathcal{B}$ is neither open nor closed.
@@
We can easily prove the complete invariance of bungee set by noting that $BU(f)=\mathbb{C}\backslash(I(f)\cup K(f))$. Now, if we assume that bungee set is closed then it will be a closed completely invariant set. This implies that $J(f)\subset BU(f)$ (because $J(f)$ is the smallest closed completely invariant set) which is a contradiction. Furthermore, if bungee set is open, it will imply that it has interior to it which is also not possible (because $BU(f)\subset J(f)$ and having interior of bungee set means Julia set also has interior which is not possible for an entire function $f$ for which $F(f)\neq\emptyset$ [3]). Hence, bungee set is neither open nor closed. ∎
@@
Functions of class $\mathcal{B}$ which are periodic translate of each other also have equal bungee sets, escaping sets and filled-in Julia sets.
@@
We take $f(z)$ and $g(z)=f(z)+C$ where $C$ is the periodic translate of $f$ (more generally, we can take $g=f^{l}(z)+C$, for some $l\in\mathbb{N}$). Then $g^{n}(z)=f^{n}(z)+C$ which is almost commuting. Here we can see that if a point is bounded under $f^{n},$ it is also bounded under $g^{n}.$ Similarly, if a point is escaping under $f^{n},$ it is also escaping under $g^{n}$. Hence, $BU(f)=BU(g),I(f)=I(g),K(f)=K(g)$. ∎
@@
Now moving forward, we will study functions that are exactly commuting of class $\mathcal{B}$ and having finite asymptotic values.
Lemma 3.4.
$g(BU(f))\subset BU(f)\cup RP(f)$ for commuting functions of $\mathcal{B}.$
Proof.
If we recall the definition of a bungee point, there must exist a subsequence which is bounded and another subsequence which is unbounded. Now it is obvious that when the subsequence which is bounded $f^{m}(z_{0})\rightarrow l_{0}$ (say) and when acted upon by $g$ will also keep it bounded (since $f,g$ are entire functions). Let us assume that there is only one such bounded subsequence and $l$ unbounded subsequence which are as mentioned below:
$f^{m}(z_{0})\rightarrow l_{0}$
$f^{n_{1}}(z_{0})\rightarrow\infty$
$f^{n_{2}}(z_{0})\rightarrow\infty$
$\vdots$
$f^{n_{l}}(z_{0})\rightarrow\infty$
Now, there are two possible cases when $g$ acts on those unbounded subsequences.
Case 1. All these subsequence gets bounded, i.e.,
$g(f^{n_{1}}(z_{0}))\rightarrow w_{1}=f^{n_{1}}(g(z_{0}))\rightarrow w_{1}$
$g(f^{n_{2}}(z_{0}))\rightarrow w_{2}=f^{n_{2}}(g(z_{0}))\rightarrow w_{2}$
$\vdots$
$g(f^{n_{l}}(z_{0}))\rightarrow w_{l}=f^{n_{l}}(g(z_{0}))\rightarrow{w_{l}}$
If this happens, then $g(z_{0})\in$ $K(f)=F(f)\cup RP(f)$. As $z_{0}\in J(f)$ and we know that $g(J(f))\subset J(f)$ hence, $g(z_{0})\in RP(f)$ for the above case. Now, let us understand what does this mean by $g(z_{0})\in RP(f)$: It means that $f^{p}(g(z_{0}))=g(z_{0})$ for some $p\in\mathbb{N}.$ This also means $g(f^{p}(z_{0}))=g(z_{0})$. Because $f$ and $g$ are not invertible functions we have $f^{p}(z_{0})\not=z_{0}$. So, $g(z_{0})\in RP(f)\ \&\ z_{0}\in BU(f)$ can be possible.
Case 2. When at least one of the subsequence still remains unbounded. Let that be the $k-$th subsequence i.e., $g(f^{n_{k}}(z_{0}))\rightarrow\infty$ i.e., $f^{n_{k}}(g(z_{0}))\rightarrow\infty$. $g(f^{m}(z_{0}))\rightarrow l_{1},$ i.e. $f^{m}(g(z_{0}))\rightarrow l_{1}$. This shows that there are at least two subsequences and in this case $g(z_{0})$ belongs to $BU(f)$. Therefore, $g(BU(f))$ is contained in either $BU(f)$ or $RP(f)$. ∎
Lemma 3.5.
$g(I(f))\subset I(f)\cup RP(f)$ for commuting functions of $\mathcal{B}.$
Proof.
Suppose $z_{0}\in I(f)$ and let $\lim_{n\rightarrow\infty}g(f^{n}(z_{0}))\rightarrow l_{0}$ (say) i.e., $f^{n}(g(z_{0}))\rightarrow l_{0}.$ This means $g(z_{0})\in K(f).$ But $z_{0}\not\in K(f)$ as it belongs to $I(f)$. So $g(z_{0})\in RP(f)$ for the above case. Now, suppose $\lim_{n\rightarrow\infty}g(f^{n}(z_{0}))\rightarrow\infty$ i.e., $f^{n}(g(z_{0}))\rightarrow\infty$. Hence, $g$ sends $I(f)$ to either $RP(f)$ or $I(f)$. So, $g(I(f))\subset I(f)\cup RP(f)$. ∎
Proposition 3.6.
$BU(f\circ g)\subset BU(f)\cup BU(g)$ for commuting functions of class $\mathcal{B}$.
Proof.
Let $z_{0}\in BU(f\circ g)\ \backslash(BU(f)\cup BU(g))$. This implies that $z_{0}\in K(f)\cup K(g)$ or $z_{0}\in I(f)\cup I(g)$. By definition of bungee set, there are subsequences $\{n_{k}\}$ and $\{m_{k}\}$ tending to $\infty$ satisfying
$|f^{n_{k}}(g^{n_{k}}(z_{0}))|\rightarrow\infty$
$|f^{m_{k}}(g^{m_{k}}(z_{0}))|\rightarrow l_{0}\ \ \text{for some }l_{0}.$
First, we assume that $z_{0}\in K(g)$ which in turn implies $z_{0}\in RP(g)$ (there is a clear contradiction in the $F(g)$ case because $F(g)=F(f\circ g)$), and as a result, $f^{n}(RP(g))\subset RP(g)$ and $g^{n}(f^{n}(RP(g)))$ which will be bounded. As can be seen, there are no subsequences that are unbounded. Similarly, there are no such points in $K(f)$. Therefore, $z_{0}\not\in K(f)\cup K(g)$.
Now, we assume $z_{0}\in I(g)$ which means there are two possible cases
(3.1) $|f(g^{n}(z_{0}))|\rightarrow l\ \text{for some}\ l$
(3.2) $|f(g^{n}(z_{0}))|\rightarrow\infty$
where $n$ denotes an arbitrary index unrelated to $n_{k}$ and $m_{k}$. If we have case (3.2), then from the array of sequences argument, we see that all subsequences of $(f\circ g)^{n}(z_{0})$ tend to infinity. If we have case (3.1), there are further possible cases: $l\in I(f)$ or $l\in K(f)$ or $l\in BU(f)$. Again, if we have the first two cases, then we observe that entire composite sequence $(f\circ g)^{n}$ is (un)bounded. The third case is also not possible. This is because using the previous Lemma (3.4), we have $|f^{n_{k}}(g^{n_{k}})(z_{0})|$ or $|f^{m_{k}}(g^{m_{k}})(z_{0})|$ which equals $|g^{n_{k}}(f^{n_{k}})(z_{0})|$ or $|g^{m_{k}}(f^{m_{k}})(z_{0})|$ will be either completely bounded or unbounded. Therefore, $z_{0}\not\in I(g)$. On similar lines, $z_{0}\not\in I(f)$. Therefore, $z_{0}\not\in I(f)\cup I(g)$. Thus, we see that there are no points s.t., $z_{0}\in BU(f\circ g)\ \backslash(BU(f)\cup BU(g))$. ∎
Corollary 3.7.
For commuting functions of class $\mathcal{B},(I(f)\cup I(g))\cap BU(f\circ g)=\emptyset$
Proof.
We can follow the similar lines of proof that is done in the preceeding Proposition (3.6). Let $z_{0}\in I(f)$ (say). Then $g(f^{n}(z_{0}))\rightarrow\infty$ or $L$. If $g(f^{n}(z_{0}))\rightarrow\infty,$ then by array of sequences argument, we conclude that $g^{n}\circ f^{n}(z_{0})\not\rightarrow M$ for some $M$ which means that we do not have a subsequence that remains bounded upon iteration of $f\circ g$. Moreover, if $g(f^{n}(z_{0}))\rightarrow L$ then $g(z_{0})\in RP(f)$ and therefore, $f^{n}(g^{n}(f^{n}(z_{0})))\rightarrow L_{1}$ (say). Thus, we never get a subsequence that tends to $\infty$.∞ . Similarly, it can be argued when $z_{0}\in I(g)$. ∎
Revision 3.8.
Revised proof of $I(f\circ g)\subset I(f)\cup I(g)$ for commuting functions of $\mathcal{B}.$ (considering the existence of bungee set)
Proof.
The motivation for the proof comes from [14]. Let $z_{0}\in I(f\circ g)\ \backslash(I(f)\cup I(g))$ i.e., $z_{0}\in RP(f)\cup RP(g)$ or $z_{0}\in BU(f)\cup BU(g)$. By definition of escaping set: $f^{n}\circ g^{n}(z_{0})=(f\circ g)^{n}(z_{0})\rightarrow\infty$. Assume $z_{0}\in BU(f)$, then $g(BU(f))$ goes to either $BU(f)$ or $RP(f)$ using Lemma 3.4. When it goes to $RP(f)$, we know $g(RP(f))$ will remain in $RP(f)$ and upon further iteration of $f\circ g$ will stay bounded i.e., $(g\circ f)^{n}(z_{0})\rightarrow L$ as $n\rightarrow\infty$ which is a contradiction. If it goes to $BU(f)$ then again, its the same case as discussed before when $g$ acts on it. In the end, it will either be an entirely bounded sequence or there would be an (un)bounded subsequence. The later case arises only when $g^{n}(BU(f))\subset BU(f)\ \forall n\in\mathbb{N}$ (where $n$ is some arbitrary index). But by definition, escaping sets must have all the subsequences tending to infinity. Hence, we arrive at a contradiction. On similar lines, it can be shown that when $z_{0}\in BU(g)$ we again arrive at a contradiction. So, $z_{0}\not\in BU(f)\cup BU(g)$.
Now the case where $z_{0}\in RP(f)\cup RP(g)$. Assuming $z_{0}\in RP(f)$ and as discussed earlier we will have all bounded subsequences of $f^{n}\circ g^{n}(z_{0})$ which is a contradiction. Similar arguments can be given when $z_{0}\in RP(g).$ Therefore $z_{0}\not\in RP(f)\cup RP(g).$ Hence, $z_{0}\in I(f\circ g)\ \backslash(I(f)\cup I(g))$ is not possible. Thus, $I(f\circ g)\subset I(f)\cup I(g)$. ∎
Proposition 3.9.
$RP(f\circ g)\supset RP(f)\cap RP(g)$ for commuting functions of class $\mathcal{B}$.
Proof.
$BU(f\circ g)\cup I(f\circ g)\subset BU(f)\cup BU(g)\cup I(f)\cup I(g)$
Taking complement on both sides
$BU(f\circ g)^{c}\cap I(f\circ g)^{c}\supset BU(f)^{c}\cap BU(g)^{c}\cap I(f)^{c}\cap I(g)^{c}$
(3.3) $\implies RP(f\circ g)\supset RP(f)\cap RP(g)$
Remark 3.10.
For $f\in\mathcal{B},$ we have $J(f)=\overline{I(f)}=\overline{BU(f)}=\overline{RP(f)}=I(f)\cup BU(f)\cup RP(f).$
@@ 4. Illustrations:
Consider the functions $f(z)=z^{2}$ and $g=1/z^{2}$. Observe that $f\circ g(z)=1/z^{4}$. $BU(f)=\emptyset$, $BU(g)=\{z:|z|<1\cup|z|>1\}$ and $BU(f\circ g)=\{z:|z|<1\cup|z|>1\}.$ Clearly, $BU(f\circ g)\subset BU(f)\cup BU(g)$.
Our next example is motivated from [15]. Let us take the case where $f=p(z)\cdot e^{h(z)}$ where $p(z),\ h(z)$ are polynomials and $g(z)=a\cdot f(z)$ for some $a\in\mathbb{C}\setminus\{0\}$. For simplicity, let us take $f(z)=z\cdot e^{a^{(k+1)}\cdot z^{k}},\ k\in\mathbb{N}$ $g(z)=a\cdot z\cdot e^{a^{(k+1)}\cdot z^{k}}$.
@@
Now, for the simplest case let us take: $k+1=2\implies a=\pm 1$ Since a=1 is a trivial case so let us take $a=-1$, $f(z)=ze^{-z^{2}};\ g(z)=-ze^{-z^{2}}$. Here as the orbit of any point under either of the function tends to $\infty,f(z)\rightarrow 0,g(z)\rightarrow 0$ and $z=0$ is a fixed point of $f(z),\ g(z)$ since $f(0)=0$ and $g(0)=0$. Now for the multiplier $\lambda$, $f^{\prime}(z)=e^{-z^{2}}(1-2z^{2}),f^{\prime}(0)=1$ which is an indifferent fixed point. And $11^{1}=$ 1 means $0\in J(f)$ or $J(g)$ (since both of them are equal). As $\infty$ is a special case and belongs to $J(f)$ so, this proves that a sequence which tends to $\infty$ if it asymptotes to a finite asymptotic value which is a member of the Julia set of $f$. Moreover, the functions which are of the form $f=p(z)e^{h(z)}$ have a finite asymptotic value ${}^{\prime}0^{\prime}$ and it is in $J(f)$ (rationally indifferent fixed point or repelling periodic point).
@@
Let $f(z)=ze^{z^{2}};\ g(z)=-ze^{z^{2}}$ $\implies f^{2}=ze^{z^{2}}e^{z^{2}e^{2z^{2}}}\ \&\ g^{2}=ze^{z^{2}}e^{z^{2}e^{2z^{2}}}.$ $f\circ g=-ze^{z^{2}}e^{z^{2}e^{2z^{2}}}=g\circ f$. And $(f\circ g)^{2}=g^{4}=f^{4}$(because as $(f\circ g)^{2}$ is done it does away with the negative term and rest terms are exactly same). So, $BU(f^{4})=BU(g^{4})=BU((f\circ g)^{2})\implies BU(f)=BU(g)=BU(f\circ g)$ (NOTE: If we have $f(z)=ze^{-z^{2}};\ g(z)=-ze^{-z^{2}}$, they also commute). We may also take $f(z)=ze^{a^{(k+1)}z^{k}}$, $g(z)=aze^{a^{(k+1)}z^{k}}$ and $a^{(k+1)}=1$ for that matter). And this complies with the union relation with the composition.
Now, for the general case: Since $a$ is $k$-th root of unity, means $a^{k}=1$. So, $f(z)=ze^{a^{k}z^{k}}=ze^{z^{k}}$ and $g(z)=aze^{z^{k}}$. Now $g^{k}(z)=a^{k}\cdot f^{k}(z)=f^{k}(z)\implies BU(f)=BU(g)$. Also, $(f\circ g)^{k}(z)=a^{k}f^{2k}(z)=f^{2k}(z)$. So, $BU(f\circ g)=BU(f)=BU(g)$. Analogously, if we take $f(z)=ze^{a^{(k+1)}z^{k}}$ and $\ g(z)=af(z)=aze^{a^{(k+1)}z^{k}}$, where again $a^{k+1}=1$, we can simplify it further as $f(z)=ze^{az^{k}}$ and $g(z)=aze^{az^{k}}$, now $g^{2}(z)=af(af(z)))=a^{2}f^{2}(z)$. By induction, $g^{k+1}(z)=a^{(k+1)}f^{k+1}(z)$. Also, $f\circ g(z)=af^{2}(z)=g\circ f(z)$ and $(f\circ g)^{2}(z)=a^{2}f^{4}(z)$. Proceeding by induction, $(f\circ g)^{k+1}(z)=a^{(k+1)}f^{2(k+1)}(z)=f^{2(k+1)}(z)$. Here also, we can see that $BU(f\circ g)=BU(f)=BU(g)$.
@@
5. Conclusion:
We have shown that entire functions with no finite asymptotic values which are also commuting i.e., $f\circ g=g\circ f$ satisfies:
$K(f)=K(g),\ I(f)=I(g),\ BU(f)=BU(g)$ and therefore, $K(f\circ g)=K(f),\ BU(f\circ g)=BU(f),\ I(f\circ g)=I(f)$.
And for the functions of class $\mathcal{B}$ which are commuting
$RP(f\circ g)\supset RP(f)\cap RP(g)$;
$BU(f\circ g)\subset BU(f)\cup BU(g)$;
$I(f\circ g)\subset I(f)\cup I(g)$.
There are some other special cases where we have equality among these sets, like when $f$ is an exponential map. Also, cases like $g=f+C$ where $C\in\mathbb{C}$ is the period of $f$. Also, when functions are from class $\mathcal{B}$, $(BU(f)\cup BU(g))\cap RP(f\circ g)\neq\emptyset$ because $g(BU(f))\subset RP(f)\cup BU(f)$. Similarly, $(I(f)\cup I(g))\cap RP(f\circ g)\neq\emptyset$ because $g(I(f))\subset RP(f)\cup I(f)$. Also, there are no points which are in $I(f)\cup I(g)$ as well as in $BU(f\circ g)$.

### v4

@@ 3. Some results for functions of class $\mathcal{B}$
Now, we will prove some results when the functions belong to the Eremenko-Lyubich class $\mathcal{B}$. These functions do not have a Baker domain; therefore, the escaping set does not intersect with the Fatou set. In addition, the bungee set of the function does not intersect with a wandering domain in which the iterates tend to $\infty.$ However, functions in this class may have oscillating wandering domain.
The next result gives a sufficient condition under which the bungee set is contained in the Julia set.
@@
The bungee set for functions of class $\mathcal{B}$ having no oscillatory wandering domain does not intersect with Fatou set, that is, $BU(f)\subset J(f).$
@@
For a function $f,$ we know that if $BU(f)\cap F(f)\not$$\emptyset$, then such components are oscillatory wandering domains [16] and these components are contained inside $BU(f)$. As $F(f)$ has no oscillatory wandering domains, it means, $BU(f)$ does not intersect with $F(f)$. In other words, $BU(f)\subset J(f).$
Remark 3.2.
This also means that $BU(f)\subset J(f)$ for functions not necessarily of $\mathcal{B}$ having no oscillatory wandering domain. Also, it is easy to see that $BU(f)$ is completely invariant by the complete invariance of $I(f)$ and $K(f)$. As a result, by the minimality property of Julia set, $\overline{BU(f)}=J(f)$. It is also known that for a function of $\mathcal{B},J(f)=\overline{I(f)}$ [6].
@@
The bungee set for functions of class $\mathcal{B}$ having no oscillatory wandering domain is neither open nor closed.
@@
We can easily prove the complete invariance of bungee set by noting that $BU(f)=\mathbb{C}\backslash(I(f)\cup K(f))$. Now, if we assume that bungee set is closed then it will be a closed completely invariant set. This implies that $J(f)\subset BU(f)$ (because $J(f)$ is the smallest closed completely invariant set) which is a contradiction as $J(f)$ contains periodic points. Furthermore, if bungee set is open, it will imply that it has interior points which is also not possible (because $BU(f)\subset J(f)$ and bungee set having interior means Julia set also has interior which is not possible for an entire function $f$ for which $F(f)\neq\emptyset$ [3]). Hence, bungee set is neither open nor closed. ∎
The following result provides equality of bungee set, filled-in Julia set and escaping set of two functions in which one function is periodic translate of the other. Note that the two functions are non-commuting.
@@
Functions which are periodic translate of each other have equal bungee sets, escaping sets and filled-in Julia sets.
@@
We consider two functions $f$ and $g$ where $g$ satisfies $g(z)=f(z)+C.$ Here, $C$ is the period of $f$ (more generally, we can take $g=f^{l}(z)+C$, for some $l\in\mathbb{N}$). It can be easily seen that $g^{n}(z)=f^{n}(z)+C$. Again, it is easy to show that if a point is bounded under $f^{n},$ it is also bounded under $g^{n}.$ Similarly, if a point is escaping under $f^{n},$ it is also escaping under $g^{n}$. Hence, $K(f)=K(g),I(f)=I(g)$ and $BU(f)=BU(g)$. ∎
@@ 4. Illustrations:
Consider the functions $f(z)=z^{2}$ and $g(z)=1/z^{2}$. Observe that $f\circ g(z)=1/z^{4}$. $BU(f)=\emptyset$, $BU(g)=\{z:|z|<1\cup|z|>1\}$ and $BU(f\circ g)=\{z:|z|<1\cup|z|>1\}.$ Clearly, $BU(f\circ g)\subset BU(f)\cup BU(g)$.
Our next example is motivated from [15]. Let us take the case where $f=p(z)\cdot e^{h(z)}$ where p(z),h(z)p(z),\ h(z) are polynomials and $g(z)=a\cdot f(z)$ for some $a\in\mathbb{C}\setminus\{0\}$. For simplicity, let us take $f(z)=z\cdot e^{a^{(k+1)}\cdot z^{k}},\ k\in\mathbb{N}$ & $g(z)=a\cdot z\cdot e^{a^{(k+1)}\cdot z^{k}}$.
@@
Now, for the simplest case let us take: $k+1=2\implies a=\pm 1$ Since $a=1$ is a trivial case so let us take a=1a=-1, $f(z)=ze^{-z^{2}};\ g(z)=-ze^{-z^{2}}$. Observe that $z=0$ is an asymptotic value of $f$ and $g$. It is also a fixed point of f(z),g(z)f(z),\ g(z) since $f(0)=0$ and $g(0)=0$. Now for the multiplier $\lambda$, $f^{\prime}(z)=e^{-z^{2}}(1-2z^{2}),f^{\prime}(0)=1$ which is an indifferent fixed point. And $1^{1}=1$ means $0\in J(f)$ or $J(g)$ (since both of them are equal). As $\infty$ is a special case and belongs to $J(f)$ so, this proves that a sequence which tends to $\infty$ if it asymptotes to a finite asymptotic value which is a member of the Julia set of $f$.
@@
Let $f(z)=ze^{z^{2}};\ g(z)=-ze^{z^{2}}.$ This implies $f^{2}=ze^{z^{2}}e^{z^{2}e^{2z^{2}}}\ \&\ g^{2}=ze^{z^{2}}e^{z^{2}e^{2z^{2}}}.$ $f\circ g=-ze^{z^{2}}e^{z^{2}e^{2z^{2}}}=g\circ f$. And $(f\circ g)^{2}=g^{4}=f^{4}$(because as $(f\circ g)^{2}$ is done it does away with the negative term and rest terms are exactly same). So, $BU(f^{4})=BU(g^{4})=BU((f\circ g)^{2})$ which implies $BU(f)=BU(g)=BU(f\circ g)$ (NOTE: If we have $f(z)=ze^{-z^{2}};\ g(z)=-ze^{-z^{2}}$, they also commute). We may also take $f(z)=ze^{a^{(k+1)}z^{k}}$, $g(z)=aze^{a^{(k+1)}z^{k}}$ and $a^{(k+1)}=1$ for that matter). And this complies with the union relation with the composition.
Now, for the general case: Since $a$ is $k$-th root of unity, means $a^{k}=1$. So, $f(z)=ze^{a^{k}z^{k}}=ze^{z^{k}}$ and $g(z)=aze^{z^{k}}$. Now $g^{k}(z)=a^{k}f^{k}(z)=f^{k}(z)$ implies $BU(f)=BU(g)$. Also, $(f\circ g)^{k}(z)=a^{k}f^{2k}(z)=f^{2k}(z)$. So, $BU(f\circ g)=BU(f)=BU(g)$. Analogously, if we take $f(z)=ze^{a^{(k+1)}z^{k}}$ and $g(z)=af(z)=aze^{a^{(k+1)}z^{k}}$, where again $a^{k+1}=1$, we can simplify it further as $f(z)=ze^{az^{k}}$ and $g(z)=aze^{az^{k}}$, now $g^{2}(z)=af(af(z)))=a^{2}f^{2}(z)$. By induction, $g^{k+1}(z)=a^{(k+1)}f^{k+1}(z)$. Also, $f\circ g(z)=af^{2}(z)=g\circ f(z)$ and $(f\circ g)^{2}(z)=a^{2}f^{4}(z)$. Proceeding by induction, $(f\circ g)^{k+1}(z)=a^{(k+1)}f^{2(k+1)}(z)=f^{2(k+1)}(z)$. Here also, we can see that $BU(f\circ g)=BU(f)=BU(g)$.